OpenStudy (anonymous):

solve by substitution method 4x+2y=-20, x=25-8y

OpenStudy (anonymous):

4x+2y=-20 Plug in 25-8y for x in the original equation 4(25-8y)+2y=-20 get rid of the parenthesis 100-32y+2y=-20 combine like terms 100-30y=-20 subtract 100 from both sides -30y=-120 divide each side by -30 to get y alone y=4

OpenStudy (anonymous):

Now that you know what y is go ahead and go back to the original problem and plug in 4 for y to find what x is.

OpenStudy (anonymous):

4x+2y=-20 4x+2(4)=-20 4x+8=-20 4x=-28 x=-7 x=-7 and y=4

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yes thank you, i didnt know what do to after the -30 thank you

OpenStudy (anonymous):

No problem just wanted to make sure you understood it:) Let me know if you need anymore help! :)